# On the Use of Agentic Coding: An Empirical Study of Pull Requests on GitHub

### v2

@@ 5.2. Implications for developers
Split large, multi-purpose PRs into smaller, easy-to-review submissions. Over 27% of Agentic-PRs combined multiple tasks (Section 4.1), and “too large” was among the top three rejection reasons (Section 4.2). Large PRs increase reviewer cognitive load and raise the risk of rejection, as previous studies (DBLP:conf/msr/WeissgerberND08) have warned developers for decades. This challenge remains relevant even for agents. In fact, RQ1 shows agents generate comprehensive solutions that attempt to address multiple issues simultaneously. The tendency of AI coding agents to produce extensive changes when given broad directives mirrors the problematic patterns seen in human development, but with potentially greater scale and complexity. When assigning a task to a coding agent, developers should adhere to the principle of providing a small, self-contained task that can be addressed in a single PR, or structure broader tasks to produce a sequence of smaller, follow-up PRs for wide-reaching changes.
Embed project-specific style, rules, and architecture in agent instructions. Style mismatches accounted for 22.122.1% of revisions, while refactors accounted for another 25.725.7% (Section 4.4). Missing documentation and tests contributed 21.7% and 13.7%, respectively. These statistics reveal a critical pattern: much of the revision work on AI-generated code results not from functional errors but from integration friction, which is the gap between what agents produce and what teams expect. This misalignment creates substantial hidden costs, as developers must repeatedly guide agents toward project-specific conventions that could have been specified upfront. We recommend that developers maintain a dedicated guideline file (e.g., CLAUDE.md, supported by Claude Code) that includes formatting rules, design principles, and architectural constraints. This practice can help coding agents produce outputs aligned with project standards, ensuring that code, documentation, and tests remain synchronized as part of the definition of done.
@@ 5.3. Implications for coding agent builders
Coding agent builders should integrate additional tools to improve adherence to project conventions. The high rates of style (22.122.1%) and refactoring (25.725.7%) revisions indicate that generic models do not fully capture local conventions (Section 4.4). This also implies that creating a persistent friction between agent capabilities and team expectations. This misalignment stems from agents being trained on diverse codebases with conflicting conventions, making them unable to infer project-specific patterns from limited context. Using the Model Context Protocol (MCP) servers (hasan2025mcp), builders can connect agents to linters (e.g., pylint) and static analysis tools to enforce formatting and structural rules before submission. Similarly, build-checking and test-coverage tools could be integrated to catch common revision triggers in advance.
@@ 6.1. Human-AI collaboration in software engineering
Recent research has emphasized the increasing collaboration between human developers and AI-powered technologies in common software engineering tasks (DBLP:journals/csur/LambiaseCPF25), such as requirements documentation (DBLP:conf/icsm/RahmanZMRRS24), code search (DBLP:conf/sigsoft/LiuZ0WH024), and code generation (DBLP:journals/jss/DakhelMNKDJ23). Hassan et al. (DBLP:journals/corr/abs-2410-06107) characterized this collaborative relationship within the emerging paradigm of Software Engineering 3.0, envisioning AI teammates that enhance developer productivity and reduce the cognitive load on human developers. However, despite these technical advances, human oversight remains crucial. AI-powered tools have demonstrated potential for improving productivity, but human-in-the-loop approaches are still required to maintain high-quality results (DBLP:conf/aiware/CoutinhoMSDFS24; DBLP:conf/xpu/MockMR24). Hence, establishing clear guidelines and best practices for managing human-AI interactions remains an essential step for industrial organizations to integrate these technologies (DBLP:conf/icst/SantosSMS24).
In addition, multi-agent frameworks such as LLM-Agent, FlowGen and FightFire assign specialised roles (planner, coder, tester) to cooperating LLMs, emulating Agile practices and raising robustness (DBLP:journals/tosem/HeTL25; DBLP:conf/icse/LinKC25; DBLP:journals/tse/YuLHKLX24). A recent systematic review consolidates prompt-engineering techniques and taxonomies of human-AI collaboration scenarios in software engineering (DBLP:journals/tosem/HouZLYWLLLGW24).
However, despite these technical achievements, significant social and human challenges exist in AI-assisted development. Piorkowski et al. (DBLP:journals/pacmhci/PiorkowskiPWWMP21) and Nahar et al. (DBLP:conf/icse/NaharZLZK25) uncover persistent knowledge gaps inside multi-disciplinary AI teams, reinforcing the call for explicit interaction guidelines. Furthermore, Adam et al. (10.1145/3721127) found that LLM-assisted reviews can disrupt collective accountability mechanisms inherent in traditional peer reviews, as developers value intrinsic factors such as professional integrity and reputation that cannot be maintained when interacting with AI systems. These studies highlight the complex socio-technical challenges in human-AI collaboration.
@@ 6.2. AI-assisted issue reports and PRs
Previous studies have explored AI-assisted activities related to issue reports (DBLP:conf/kbse/BoJ000024; DBLP:conf/icse/KangYY23) and PRs (DBLP:conf/msr/ChouchenBBOAM24; DBLP:conf/ease/WatanabeK0HYI24). However, no prior research has investigated the use of agentic coding tools, such as OpenAI’s Codex [^47] https://openai.com/index/introducing-codex and Claude Code, in real-world development contexts. Instead, current research primarily focuses on LLM-based chatbots like ChatGPT. For example, Chouchen et al. (DBLP:conf/msr/ChouchenBBOAM24) reported that ChatGPT is frequently used in review-intensive PRs to reduce developer effort. Despite this positive impact, 30.7% of AI-generated suggestions during code reviews were still met with skepticism or disagreement (DBLP:conf/ease/WatanabeK0HYI24). Our study fills this gap by tracing 567 autonomous PRs created by Claude Code across 157 open-source projects, reporting acceptance rates, revision effort, and rejection rationales.
Researchers have proposed various approaches for directly involving LLMs in activities related to issue reports and PRs. For example, Kang et al. (DBLP:conf/icse/KangYY23) used LLMs to generate test cases from issue reports for easier bug reproduction, Zhang et al. (DBLP:conf/icsm/IrsanZT0J22) trained LLMs to automatically generate PR titles, and Bo et al. (DBLP:conf/kbse/BoJ000024) leveraged ChatGPT to improve the clarity of issue reports for helping subsequent tasks such as bug fixing.
Complementary work investigates AI agents that clarify ambiguous requirements and collaboratively co-evolve PR descriptions; for instance, AgileGen aligns end-user perspectives with acceptance criteria (DBLP:journals/tosem/ZhangXGXCZZFZ25), while large-scale analyses of LLMs on code-change tasks reveal their potential for automated review of PR diffs (DBLP:journals/tosem/FanLLLXL25). Multi-agent benchmarks also examine ChatGPT’s ability to self-verify its generated code through test-report generation, reducing reviewer burden (DBLP:journals/tse/YuLHKLX24).
Early work by Tufano et al. (DBLP:conf/icse/TufanoPTPB21) pioneered DL models that propose code changes before review submission and apply reviewer suggestions after submission. Guo et al. report that GPT-4 can refine 66% of files given review comments (DBLP:conf/icse/GuoCXLL0024), and later boost intent-extraction accuracy to 79% in their ICSE ’25 study (DBLP:conf/icse/GuoXL00B25). To improve developer confidence, Di and Zhang (DBLP:conf/icse/Di025) interleave inline comments with LLM edits, cutting task time by 16.7 percent.
@@ 6.3. Empirical studies of automatic code generation
Researchers have studied various aspects of AI-generated code, including correctness (DBLP:conf/esem/BillahRCR24), efficiency (DBLP:conf/forge/NiuZL0024), and security (DBLP:journals/tse/LiuTLZZ24). For example, Billah et al. (DBLP:conf/esem/BillahRCR24) evaluated the correctness (pass rate) of LLMs in solving programming challenges on competitive programming platforms such as LeetCode and Codeforces. They reported a decrease in the pass rate in live Codeforces contests compared to archived problems. However, code correctness alone does not necessarily imply efficiency or optimality as code complexity increases, and extra effort is required to support AI models in generating efficient code (DBLP:conf/forge/NiuZL0024). From a security perspective, Liu et al. (DBLP:journals/tse/LiuTLZZ24) identified security vulnerabilities in AI-generated code and highlighted potential security risks that must be mitigated before deployment in production environments. Beyond generating code at the granularity of individual functions or files, SWE-bench (DBLP:conf/iclr/JimenezYWYPPN24) was introduced to evaluate the performance of AI-powered code generation tools in resolving real-world GitHub issues.
Subsequent work has shown that ChatGPT often produces buggy or low-quality code that benefits from iterative repair (DBLP:journals/tse/LiuTLZZ24), while self-collaboration frameworks with multiple LLM agents significantly improve correctness and reduce human intervention (DBLP:journals/tosem/DongJJL24). Empirical studies highlight the challenges of domain-specific coding (e.g., machine-learning pipelines) and demonstrate that knowledge-prompting strategies boost performance in such settings (DBLP:journals/tosem/GuCLHZWWXW25; DBLP:journals/tosem/ShinWWSW24). Recent surveys map out future research directions for automatic programming pipelines, including integrated quality-assurance stages (DBLP:journals/tosem/LyuRRTT25), and bias-testing frameworks reveal that LLMs can inherit societal biases, motivating prompt-engineering mitigation strategies (10.1145/3724117).
Wang et al. (DBLP:conf/icse/WangZSHCMZ25) catalogue 557 distinct error patterns that six LLMs make on HumanEval. Niu et al. benchmark 19 pre-trained models across 13 tasks, confirming that functional correctness rarely correlates with efficiency (DBLP:conf/icse/NiuLNCGL23). O’Brien et al. (DBLP:conf/icse/OBrienBIASR24) show Copilot frequently injects TODO-style self-admitted technical debt. For repair, Xia et al. (DBLP:conf/icse/XiaWZ23) and Bouzenia et al. (DBLP:conf/icse/BouzeniaDP25) re-cast bug fixing as an autonomous planning problem executed by an LLM agent.

### v3

@@ 5.2. Implications for developers
Split large, multi-purpose PRs into smaller, easy-to-review submissions. Nearly 40% of Agentic-PRs combined multiple tasks (Section 4.1), and “too large” was among the top three rejection reasons (Section 4.2). Large PRs increase reviewer cognitive load and raise the risk of rejection, as previous studies (Weißgerber et al., 2008) have warned developers for decades. This challenge remains relevant even for agents. In fact, RQ1 shows agents generate comprehensive solutions that attempt to address multiple issues simultaneously. The tendency of AI coding agents to produce extensive changes when given broad directives mirrors the problematic patterns seen in human development, but with potentially greater scale and complexity. When assigning a task to a coding agent, developers should adhere to the principle of providing a small, self-contained task that can be addressed in a single PR, or structure broader tasks to produce a sequence of smaller, follow-up PRs for wide-reaching changes.
Embed project-specific style, rules, and architecture in agent instructions. Style mismatches accounted for 23.423.4% of revisions, while refactors accounted for another 27.127.1% (Section 4.4). Missing documentation and tests contributed 29.029.0% and 16.416.4%, respectively. These statistics reveal a critical pattern: much of the revision work on AI-generated code results not from functional errors but from integration friction, which is the gap between what agents produce and what teams expect. This misalignment creates substantial hidden costs, as developers must repeatedly guide agents toward project-specific conventions that could have been specified upfront. We recommend that developers maintain a dedicated guideline file (e.g., CLAUDE.md, supported by Claude Code) that includes formatting rules, design principles, and architectural constraints. This practice can help coding agents produce outputs aligned with project standards, ensuring that code, documentation, and tests remain synchronized as part of the definition of done.
@@ 5.3. Implications for coding agent builders
Coding agent builders should integrate additional tools to improve adherence to project conventions. The high rates of style (23.423.4%) and refactoring (27.127.1%) revisions indicate that generic models do not fully capture local conventions (Section 4.4). This also implies that it creates persistent friction between agent capabilities and team expectations. This misalignment stems from agents being trained on diverse codebases with conflicting conventions, making them unable to infer project-specific patterns from limited context. Using the Model Context Protocol (MCP) servers (Hasan et al., 2025), builders can connect agents to linters (e.g., pylint) and static analysis tools to enforce formatting and structural rules before submission. Similarly, build-checking and test-coverage tools could be integrated to catch common revision triggers in advance.
@@ 6.1. Human-AI collaboration in software engineering
Recent research has emphasized the increasing collaboration between human developers and AI-powered technologies in common software engineering tasks (Lambiase et al., 2025), such as requirements documentation (Rahman et al., 2024), code search (Liu et al., 2024a), and code generation (Dakhel et al., 2023). Hassan et al. (Hassan et al., 2024) characterized this collaborative relationship within the emerging paradigm of Software Engineering 3.0, envisioning AI teammates that enhance developer productivity and reduce the cognitive load on human developers. However, despite these technical advances, human oversight remains crucial. AI-powered tools have demonstrated potential for improving productivity, but human-in-the-loop approaches are still required to maintain high-quality results (Coutinho et al., 2024; Mock et al., 2024). Hence, establishing clear guidelines and best practices for managing human-AI interactions remains an essential step for industrial organizations to integrate these technologies (Santos et al., 2024).
In addition, multi-agent frameworks such as LLM-Agent, FlowGen and FightFire assign specialised roles (planner, coder, tester) to cooperating LLMs, emulating Agile practices and raising robustness (He et al., 2025; Lin et al., 2025; Yu et al., 2024). A recent systematic review consolidates prompt-engineering techniques and taxonomies of human-AI collaboration scenarios in software engineering (Hou et al., 2024).
However, despite these technical achievements, significant social and human challenges exist in AI-assisted development. Piorkowski et al. (Piorkowski et al., 2021) and Nahar et al. (Nahar et al., 2025) uncover persistent knowledge gaps inside multi-disciplinary AI teams, reinforcing the call for explicit interaction guidelines. Furthermore, Adam et al. (Alami et al., 2025) found that LLM-assisted reviews can disrupt collective accountability mechanisms inherent in traditional peer reviews, as developers value intrinsic factors such as professional integrity and reputation that cannot be maintained when interacting with AI systems. These studies highlight the complex socio-technical challenges in human-AI collaboration.
@@ 6.2. AI-assisted issue reports and PRs
Previous studies have explored AI-assisted activities related to issue reports (Bo et al., 2024; Kang et al., 2023) and PRs (Chouchen et al., 2024; Watanabe et al., 2024). However, no prior research has investigated the use of agentic coding tools, such as OpenAI’s Codex [^47] https://openai.com/index/introducing-codex and Claude Code, in real-world development contexts. Instead, current research primarily focuses on LLM-based chatbots like ChatGPT. For example, Chouchen et al. (Chouchen et al., 2024) reported that ChatGPT is frequently used in review-intensive PRs to reduce developer effort. Despite this positive impact, 30.7% of AI-generated suggestions during code reviews were still met with skepticism or disagreement (Watanabe et al., 2024). Our study fills this gap by tracing 567 autonomous PRs created by Claude Code across 157 open-source projects, reporting acceptance rates, revision effort, and rejection rationales.
Researchers have proposed various approaches for directly involving LLMs in activities related to issue reports and PRs. For example, Kang et al. (Kang et al., 2023) used LLMs to generate test cases from issue reports for easier bug reproduction, Irsan et al. (Irsan et al., 2022) trained LLMs to automatically generate PR titles, and Bo et al. (Bo et al., 2024) leveraged ChatGPT to improve the clarity of issue reports for helping subsequent tasks such as bug fixing.
Complementary work investigates AI agents that clarify ambiguous requirements and collaboratively co-evolve PR descriptions; for instance, AgileGen aligns end-user perspectives with acceptance Zhang (Zhang et al., 2025), while large-scale analyses of LLMs on code-change tasks reveal their potential for automated review of PR diffs (Fan et al., 2025). Multi-agent benchmarks also examine ChatGPT’s ability to self-verify its generated code through test-report generation, reducing reviewer burden (Yu et al., 2024).
Early work by Tufano et al. (Tufano et al., 2021) pioneered DL models that propose code changes before review submission and apply reviewer suggestions after submission. Guo et al. report that GPT-4 can refine 66% of files given review comments (Guo et al., 2024), and later boost intent-extraction accuracy to 79% in their ICSE ’25 study (Guo et al., 2025). To improve developer confidence, Di and Zhang (Di and Zhang, 2025) interleave inline comments with LLM edits, cutting task time by 16.7 percent.
@@ 6.3. Empirical studies of automatic code generation
Researchers have studied various aspects of AI-generated code, including correctness (Billah et al., 2024), efficiency (Niu et al., 2024), and security (Liu et al., 2024b). For example, Billah et al. (Billah et al., 2024) evaluated the correctness (pass rate) of LLMs in solving programming challenges on competitive programming platforms such as LeetCode and Codeforces. They reported a decrease in the pass rate in live Codeforces contests compared to archived problems. However, code correctness alone does not necessarily imply efficiency or optimality as code complexity increases, and extra effort is required to support AI models in generating efficient code (Niu et al., 2024). From a security perspective, Liu et al. (Liu et al., 2024b) identified security vulnerabilities in AI-generated code and highlighted potential security risks that must be mitigated before deployment in production environments. Beyond generating code at the granularity of individual functions or files, SWE-bench (Jimenez et al., 2024) was introduced to evaluate the performance of AI-powered code generation tools in resolving real-world GitHub issues.
Subsequent work has shown that ChatGPT often produces buggy or low-quality code that benefits from iterative repair (Liu et al., 2024b), while self-collaboration frameworks with multiple LLM agents significantly improve correctness and reduce human intervention (Dong et al., 2024). Empirical studies highlight the challenges of domain-specific coding (e.g., machine-learning pipelines) and demonstrate that knowledge-prompting strategies boost performance in such settings (Gu et al., 2025; Shin et al., 2024). Recent surveys map out future research directions for automatic programming pipelines, including integrated quality-assurance stages (Lyu et al., 2025), and bias-testing frameworks reveal that LLMs can inherit societal biases, motivating prompt-engineering mitigation strategies (Huang et al., 2025).
Wang et al. (Wang et al., 2025) catalogue 557 distinct error patterns that six LLMs make on HumanEval. Niu et al. benchmark 19 pre-trained models across 13 tasks, confirming that functional correctness rarely correlates with efficiency (Niu et al., 2023). O’Brien et al. (O’Brien et al., 2024) show Copilot frequently injects TODO-style self-admitted technical debt. For repair, Xia et al. (Xia et al., 2023) and Bouzenia et al. (Bouzenia et al., 2025) re-cast bug fixing as an autonomous planning problem executed by an LLM agent.